COMPARATIVE STUDIES of J′ ← J = 1 ← 0 CO LINE BROADENING PARAMETERS IN FREQUENCY AND TIME DOMAIN EXPERIMENT

H.MÄDER, A.GUARNIERI, J.DOOSE, N.NISSEN, Institut für Physikalische Chemie, Universität Kiel, Olshausenstrasse 40, D-24098 Kiel, Germany; \underline{V.MARKOV}, G.GOLUBYATNIKOV, I.LEONOV, and A.F.KPUPNOV, Institute of Applied Physics of Russian Academy of Science, Ul’yanova Street 46, 603600 Niznny Novgorod, Russia.

The pressure broadening of CO J′ ← J = 1 ← 0 line has been studied by two different techniques. The investigation was made with the perturber gases CO ^{a}, N_{2}, O_{2}, H_{2}, CO_{2}, He, Ne, Ar, Kr, Xe at room temperature. The two different methods of line broadening measurement and the results obtained in the time- and in the frequencydomain respectively will be discused. Both methods gave results in a good agreement, for example, selfbroadening parameter C_{w}=3.43(2) MHz/Torr (time-domain) and C_{w}=3.425(6) (frequency-domain) were obtained. ^{b}

aH.Maeder, A.Guarnieri, J.Doose, N.Nissen, V.N.Markov, A.M.Stanyuk, A.F.Andrianov, V.N.Shanin and A.F.Krupnov, Journ.

Molec. Spectr. \underline{180}, 183-187 (1996)

bSupport by DFG and RFBR grants is gratefully acknowledged